OpenStudy (anonymous):

Parallelogram DEFG is below. DH = y + 4, HF = 3x – 1, GH = 2y, and HE = 3x + 2. Find the values of x and y. A. x = 14, y = 11 B. x = 4, y = 7 C. x = 7, y = 4 D. x = 11, y = 14

Directrix (directrix):

Where are the Hs in Parallelogram DEFG?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

The diagonals of a parallelogram bisect each other. y + 4 = 3x - 1 Also, 3x + 2 = 2y The task is to solve the pair of simultaneous equations: y + 4 = 3x - 1 3x + 2 = 2y ----------- x = 4 and y = 7 ---> Option B

OpenStudy (anonymous):

I think this one is D !

Directrix (directrix):

Let's see what I get.

Directrix (directrix):

Using (x+1 =2y) and (2x-2 = 2y+2), I got x = 5 and y = 3. ----> C

OpenStudy (anonymous):

ohh that makes since thanks !